Sam wants to swim in a river with a current of 2 miles an hour. He can travel 8 miles upstream and back in 3 hours. How fast is he in still water?
Answer by lwsshak3(11628)

let x =swim speed of of sam in still water
time = distance÷speed
downstream speed of sam =stillwater + current speed=(x+2)
upstream speed of sam= stillwater -current speed=(x-2)
time downstream + time upstream = 3
8/(x+2) +8/(x-2) = 3
LCD: (x=2)(x-2)
8(x-2)+8(x+2) =3(x+2)(x-2)=3(x^2-4)
8x-16+8x+16=3x^2-12
3x^2-16x-12 =0
(3x+2)(x-6) =0
x=6 or x=-2/3 (rejected, speed cannot be negative)
ans: swim speed of sam in still water = 6 mi/hr